OpenStudy (anonymous):

What is the volume of a cone that has a radius of 4 cm. and a height of 9 cm? Use 3.14 to approximate pi and round the answer to two decimal places.

Directrix (directrix):

V = (1/3) B*h where B is the area of the circular base and h is the height of the cone.

Directrix (directrix):

V = (1/3) B*h V = (1/3) * pi * 4^2 * 9 V = (1/3) * 3.14 * 4^2 * 9 V = @lisa2005 Post what you get, okay?

OpenStudy (anonymous):

150.79645

Directrix (directrix):

Round the answer to two decimal places. I got 150.72 sq. cm.